OpenStudy (anonymous):

Simplify 6 to the power of 6 / by 6 to the power of 5 using the laws of exponents. Verify your answer by expanding the powers, then dividing. Show your work.

OpenStudy (anonymous):

need help fast

OpenStudy (anonymous):

\[\frac{6^6}{6^5}=6^{6-5}=6^1=6\]

OpenStudy (mathstudent55):